# Manolis Mylonas CERTH-ITI Thessaloniki, Greece

### v1

@@ Abstract
In this work, we introduce the task of script-driven video summarization, which aims to produce a summary of the full-length video by selecting the parts that are most relevant to a user-provided script outlining the visual content of the desired summary. Following, we extend a recently-introduced large-scale dataset for generic video summarization (VideoXum) by producing natural language descriptions of the different human-annotated summaries that are available per video. In this way we make it compatible with the introduced task, since the available triplets of “video, summary and summary description” can be used for training a method that is able to produce different summaries for a given video, driven by the provided script about the content of each summary. Finally, we develop a new network architecture for script-driven video summarization (SD-VSum), that relies on the use of a cross-modal attention mechanism for aligning and fusing information from the visual and text modalities. Our experimental evaluations demonstrate the advanced performance of SD-VSum against state-of-the-art approaches for query-driven and generic (unimodal and multimodal) summarization from the literature, and document its capacity to produce video summaries that are adapted to each user’s needs about their content.
@@
Keywords Script-driven video summarization  $\cdot$ Dataset  $\cdot$ Natural language descriptions  $\cdot$ Cross-modal attention  $\cdot$ Large Multimodal Models
1 Introduction
@@
The established approach in the Media industry for producing summarized versions of a full-length video for distribution via different communication channels (e.g., social networks and video sharing platforms), requires the visual inspection of the entire content by a professional editor who then decides about the parts that should be included in the summary. This time-demanding and laborious task (especially in the case of long videos with diverse visual content) can be facilitated by modern AI-based video summarization technologies [1]. Nevertheless, these technologies can assist the work of editors only to a certain extent, since they typically learn how to create video summaries based on general-purpose summarization criteria, such as the representativeness and diversity of the summary (see the top part of Fig. 1), and thus, are not capable of generating video summaries according to specific needs of the editor about their content.
@@
Figure 1: A video blog (vlog) with the daily activities of the vlogger, and the produced summaries by: generic video summarization (top row), query-driven video summarization (middle row), and script-based video summarization (bottom row). The generic video summary aims to provide a complete overview of the video. The query-driven video summary concentrates solely on parts of the video that match with a short user query. The script-driven video summary offers a more diverse and complete synopsis of the video, containing parts that relate to the different activities outlined in the user script, and avoiding parts the are less interesting for the user, and thus redundant.
@@
Being more tailored to the aforementioned need, a number of text/query-driven video summarization methods have been described in the literature [2, 3, 4, 5, 6, 7, 8]. These methods allow to generate video summaries that are more customized to the users’ preferences. However, these preferences are typically expressed using one or more keywords [2, 3, 6] or a short sentence [4, 5]. Hence, the existing methods for text/query-driven video summarization are not compatible with more extensive descriptions about the content of the summary. As a consequence, the generated summaries by these methods exhibit limited visual and semantic diversity, as they mainly contain parts of the video that match a short user query (see the middle part of Fig. 1).
@@
To overcome this limitation, in this paper we introduce the task of script-driven video summarization, whose goal is to generate a summary of the full-length video by choosing the parts which are most relevant to a longer user-provided query (“script”) that outlines the visual content of the desired summary. To support research on this task, we extend a recently-introduced large-scale dataset for video summarization (VideoXum [9]), by producing natural language descriptions of the visual content of the different human-annotated summaries that are available per video, using the state-of-the-art Large Multimodal Model LLaVA-NeXT-7B [10]. The multiple triplets of “video, summary and summary description” that are available for each video in the extended S-VideoXum dataset can be used to train a method to produce different summaries for a given video, based on the provided description (the so-called script) about the content of each summary. Finally, we build a new method for script-driven video summarization (called SD-VSum), that employs a cross-modal attention mechanism for aligning and fusing information from the visual and text modalities, and computes the frames’ importance using a trainable Transformer-based scoring mechanism. SD-VSum is designed to generate video summaries that are visually and semantically more diverse (see the bottom part of Fig. 1) compared to summaries created by text/query-driven summarization methods, and less redundant compared to summaries produced by generic summarization approaches. We train and evaluate SD-VSum using S-VideoXum and compare its performance against state-of-the-art approaches for query-driven and generic (unimodal and multimodal) video summarization from the literature. Our quantitative and qualitative evaluations highlight the advanced performance of SD-VSum and demonstrate its capacity to generate video summaries that meet the users’ needs. Our contributions are as follows:
@@
  • We extend the recently-introduced VideoXum large-scale dataset for video summarization to S-VideoXum, by augmenting the former with natural language descriptions of the different human-annotated summaries that are available per video.
@@
  • We propose a new method for script-driven video summarization (SD-VSum) that employs a cross-modal attention mechanism; we use the extended S-VideoXum dataset to train and evaluate this method, and document its competitiveness against other state-of-the-art for query-driven and generic (unimodal and multimodal) video summarization.
2 Related Work
2.1 Text/Query-driven video summarization
@@
In one of the first attempts towards text-driven video summarization, Sharghi et al. (2016) [2] proposed the use of a probabilistic model for query-focused video summarization, which creates a summary by selecting the video shots that jointly have the highest relevance to the user query and the highest importance in the context of the video. To avoid the use of costly supervision data, in their subsequent work, Sharghi et al. (2017) [3] described a memory network parameterized probabilistic model that implicitly attends the user query about the video onto different video frames and shots. Vasudevan et al. (2017) [11] utilized the framework of submodular optimization to form a method for query-adaptive video summarization that selects a subset of frames for inclusion in the video summary by quantifying their relevance with the text query based on the distance of deep visual and text representations in a common embedding space, and taking into account the frames’ diversity, representativeness and aesthetic quality. Wei et al. (2018) [4] presented a semantic attended video summarization network that contains a frame selector and a video descriptor and learns how to choose the most semantically representative video parts by minimizing the distance between the generated description for the video summary, and human-provided descriptions of visually-coherent segments of the video. Zhang et al. (2018) [12] suggested the use of a query-conditioned generative adversarial network, where the generator learns the joint representation of the user query and the video content, the discriminator tries to discriminate the real summary from a generated and a random one, and a three-player loss is used to force the generator to learn the task and avoid random trivial solutions. Jiang et al. (2019) [13] designed a hierarchical network architecture to model query-related long-range temporal dependency and used diverse attention mechanisms to encode query-related and context-important information in combination with a multi-level self-attention and a variational autoencoder to incorporate user-oriented diversity and stochastic factors. Huang et al. (2020) [5] described a method for query-controllable video summarization, that consists of: a summary controller that uses text (e.g., words, phrases, sentences) to describe the desired video summary, a summary generator that scores the video frames based on their implicit relationship with the text-based description, and a summary output module that creates a video summary by selecting the top-k scoring frames. Xiao et al. (2020) [14] proposed the use of a fully convolutional network architecture which encodes the high-level semantics of each video shot using local self-attention mechanisms and models the semantic relationship between all shots and a given query using a query-aware global-attention, in combination with a query-relevance ranking module which computes a similarity score for each shot for a given query and forms the summary. Aiming to eliminate the bias when the model is trained on a small dataset, in their next work, Xiao et al. (2020) [15] trained a hierarchical self-attentive network for estimating the importance of video frames and shots using the ActivityNet Captions dataset [16], fine-tuned it using a reinforced caption generator that aims to construct the video description using the selected frames, and employed a query-aware scoring module that computes shot-level scores for a given query and produces the query-related video summary. Cizmeciler et al. (2021) [17] presented a weakly-supervised query-focused video summarization method that makes use of semantic attributes as an indicator of query relevance and semantic saliency maps obtained from predictions of pretrained attribute/action classifiers to locate relevant regions in the video frames, and employs both within a submodular maximization framework. Narasimhan et al. (2021) [6] introduced the CLIP-It method for query-focused video summarization, which relies on a multi-head language-guided attention mechanism that learns how to estimate the frames importance based on their visual relevance and their correlation with a user-defined query. Formulating the task of query-attentive video summarization as a multi-label classification problem, Hu et al. (2023) [18] predicted the correlation between visual content and multi-concept labels by feeding convolutional features into a multi-layer perceptron, used the labels’ cross-correlation to weight the predicted probability, and selected the video part with the highest relevance to the user’s query sentence to form the summary. Finally, to deal with the lack of large labeled datasets for query-based video summarization, Huang et al. (2023) [19] resorted to the use of self-supervision and the generation of segment-level pseudo labels from input videos to model the relationship between a pretext and a target task and between the pseudo label and the human-defined label, and employed a semantics booster to generate context-aware query representations and mutual attention to capture the interaction between visual and textual modalities. For further details, we refer the reader to a recently published survey on query-attentive video summarization [20].
@@
The methods described above create a video summary taking into account the user’s preferences about its content, expressed using a few keywords [2, 3, 11, 12, 13, 5, 14, 17, 15, 6, 18, 19] or a short sentence [4, 6]. Contrary to these methods, SD-VSum guides the summarization process using a free-form language description of the content of the desired video summary, allowing the user to create summaries that are visually and semantically more diverse compared to the summaries generated by the existing text/query-driven summarization methods.
2.2 Multimodal video summarization
@@
Closely related to the text/query-driven videos summarization methods, in terms of the used data modalities, are multimodal approaches for generic video summarization which take as input the full-length video and a textual description of its content and/or the video transcripts. The CLIP-It method of Narasimhan et al. (2021) [6] for query-based video summarization is also capable to produce generic video summaries using automatically-generated dense captions of the video content. In their subsequent work, Narasimhan et al. (2022) [7] described a transcript-based method for summarizing instructional videos by selecting important steps of the procedure that are most relevant to the task, as well as salient in the video, i.e., referenced in the speech. Zhong et al. (2022) [21] presented an unsupervised video summarization approach that learns how to estimate the frames’ importance and create semantically representative video summaries by minimizing the distance of learnable representations of the video content and its textual description, in a common latent space. He et al. (2023) [8] presented a unified multimodal transformer-based model which can align and attend multimodal inputs leveraging time correspondence (i.e., video and transcript), and selects the most importance frames of the video (to form a keyframe-based video summary) and the most importance sentences of the transcript (to produce a text summary). Finally, Argaw et al. (2024) [22] trained a multimodal method for video summarization using ground-truth pseudo-summaries obtained by prompting an LLM to extract the most critical and informative moments from ASR transcripts and mapping back the resulting textual summary to the relevant video segments. Their method creates a video summary by taking into account the visual content and a long-form description of it or the video transcripts.
@@
Table 1: Overview of the different datasets for text/query-driven video summarization from the literature. The original VideoXum dataset that is used as a basis for our S-VideoXum dataset, is included as a reference.
@@
UTE [23] 1 4 3 - 5 hr. 3 textual summaries train, test
TVE [24] 1 4 45 min. 3 textual summaries train, test
QFVS [3] 1 4 3 - 5 hr. 46 query-based video summaries train, test
RAD [11] 22 190 2 - 3 min. 5 query-relevance labels train, validation, test
SumMe [25] - 25 1 - 6 min. 3-5 fragment-level descriptions train, test
TVSum [26] 10 50 1 - 11 min. 3-5 fragment-level descriptions train, test
ARS [17] 2 10 2 - 12 min. 10 query-based video summaries train, test
VideoXum [9] open domain 14,000 up to 12.5 min. 10 - train, validation, test
@@ 2.3 Datasets
An overview of the datasets that were used in the literature for text/query-driven video summarization is provided in Table 1. The earliest used ones by Sharghi et al. in [2], were the UT Egocentric (UTE) [23] and the TV Episodes (TVE) [24] datasets. UTE contains four egocentric videos $3-5$ hours long that were recorded in an uncontrolled environment from a first-person view wearable camera. TV Episodes includes four videos $45$ min. long that were captured from a third person’s viewpoint. The videos of the UTE and TV Episodes have been annotated by human subjects in [24] with one-sentence descriptions at the level of short video fragments that last $5$ and $10$ sec., respectively. Three reference text-based summaries were defined per video as a subset of the textual annotations, and used for evaluating the output of the query-driven summarization method based on the ROUGE-SU metric [27]. The UTE dataset was further extended by Sharghi et al. in [3], by adding human-annotated video summaries for $46$ user queries that were defined using two or three concepts per video, forming the Query-Focused Video Summarization (QFVS) dataset. Vasudevan et al. [11] introduced the RAD dataset which contains $190$ videos that were retrieved from YouTube using different queries and annotated with query relevance labels at the frame and fragment level. Labels range from 0 (Bad) to 3 (Very Good), where “Bad” indicates that the frame is both irrelevant and of low aesthetic quality (e.g. blurred, bad contrast, etc.). This dataset was extended by Huang et al. in [5], by computing relevance scores for each query-video pair of the dataset. The SumMe [25] and TVSum [26] benchmarking datasets for generic video summarization were extended by Wei et al. [4] with fragment-level descriptions of the visual content, to train a video summarization method that takes into account the video semantics. Finally, Cizmeciler et al. [17] introduced the Activity Related Summaries (ARS) dataset that contains $10$ YouTube videos of $2$-$12$ min. duration. These videos were annotated by humans at the shot level using a pre-defined list of $60$ attributes, and $10$ ground-truth summaries were defined for each video using one/two-word queries from a list of $142$ queries.
@@
As shown in Table 1, contrary to the existing datasets, the extended S-VideoXum dataset for script-driven video summarization: i) contains a remarkably large number of videos from different domains with diverse visual content, ii) provides $10$ different ground-truth summaries per video, thus allowing to train a summarization method to produce different summaries for a given video based on the users’ preferences about the content of each summary, and iii) is divided in train, validation and test sets, enabling the selection of a well-trained model that generalizes well to unseen data.
3 Problem Statement and Dataset
3.1 Problem formulation

### v2

# SD-VSum: A Method and Dataset for Script-Driven Video Summarization

@@ Abstract.
In this work, we introduce the task of script-driven video summarization, which aims to produce a summary of the full-length video by selecting the parts that are most relevant to a user-provided script outlining the visual content of the desired summary. Following, we extend a recently-introduced large-scale dataset for generic video summarization (VideoXum) by producing natural language descriptions of the different human-annotated summaries that are available per video. In this way we make it compatible with the introduced task, since the available triplets of “video, summary and summary description” can be used for training a method that is able to produce different summaries for a given video, driven by the provided script about the content of each summary. Finally, we develop a new network architecture for script-driven video summarization (SD-VSum), that employs a cross-modal attention mechanism for aligning and fusing information from the visual and text modalities. Our experimental evaluations demonstrate the advanced performance of SD-VSum against SOTA approaches for query-driven and generic (unimodal and multimodal) summarization from the literature, and document its capacity to produce video summaries that are adapted to each user’s needs about their content.
@@
Script-driven video summarization, Dataset, Natural language descriptions, Cross-modal attention, Large Multimodal Models
journalyear: 2025copyright: ccconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755821isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Video summarizationccs: Computing methodologies Supervised learningccs: Computing methodologies Learning to rankccs: Computing methodologies Natural language generationccs: Computing methodologies Image representations
1. Introduction
@@
The established approach in the Media industry for producing summarized versions of a full-length video for distribution via different communication channels (e.g., social networks and video sharing platforms), requires the visual inspection of the entire content by a professional editor who then decides about the parts that should be included in the summary. This time-demanding and laborious task can be facilitated by modern AI-based video summarization technologies (Apostolidis et al., 2021a). Nevertheless, these technologies can assist the work of editors only to a certain extent, since they typically learn how to create video summaries based on generic summarization criteria, such as the representativeness and diversity of the summary (see the top part of Fig. 1), and thus, are not able to generate video summaries according to the editor’s needs about their content.
@@
Figure 1. A video blog (vlog) with the daily activities of the vlogger, and the produced summaries by: generic (top row), query-driven (middle row), and script-driven video summarization (bottom row). The generic video summary aims to provide a complete overview of the video. The query-driven video summary concentrates solely on parts of the video that match with a short user query. The script-driven video summary offers a more tailored synopsis of the video, containing parts that relate to the different activities outlined in the user script and avoiding parts that are less interesting for the user, and thus redundant.
@@
Being more tailored to the aforementioned requirement, a number of text/query-driven video summarization methods have appeared (Sharghi et al., 2016, 2017; Wei et al., 2018; Huang and Worring, 2020; Narasimhan et al., 2021, 2022; He et al., 2023). These methods allow to generate video summaries that are more customized to the users’ preferences. However, these preferences are typically expressed using one or more keywords (e.g., “food”, “cooking pancakes”) (Sharghi et al., 2016, 2017; Narasimhan et al., 2021) or a short sentence (e.g., “a woman is making a sandwich”) (Wei et al., 2018; Huang and Worring, 2020); as a consequence, the generated summaries by these methods show limited visual and semantic diversity, as they mainly contain parts of the video that match such a short user query (see the middle part of Fig. 1). The existing text/query-driven video summarization methods are not compatible with more extensive descriptions, formed by multiple sentences that can relate to different events depicted in the video, and possibly include abstract cues and narrative elements. This is a different problem that necessitates a different technical approach for solving it.
@@
Based on the above, in this paper we introduce the task of script-driven video summarization, whose goal is to generate a summary of the full-length video by choosing the parts which are most relevant to a longer user-provided query (“script”) that outlines the content of the desired summary. To support research on this task, we extend a recently-introduced large-scale dataset for video summarization (VideoXum (Lin et al., 2024)), by producing natural language descriptions of the visual content of the different human-annotated summaries that are available per video, using the SOTA Large Multimodal Model LLaVA-NeXT-Video-7B (Li et al., 2024). The multiple triplets of “video, summary and summary description” that are available for each video in the extended S-VideoXum dataset can be used to train a method to produce different summaries for a given video, based on the provided description (the so-called script) about the content of each summary. Finally, we build a new method for script-driven video summarization (called SD-VSum), that employs a cross-modal attention mechanism to align and fuse information from the visual and text modalities, and computes the frames’ importance using a trainable Transformer-based scoring mechanism. SD-VSum is designed to generate video summaries that are visually and semantically more diverse (see the bottom part of Fig. 1) compared to summaries created by text/query-driven summarization methods, and less redundant compared to summaries produced by generic summarization approaches. We train and evaluate SD-VSum using S-VideoXum and compare its performance against SOTA approaches for query-driven and generic (unimodal and multimodal) video summarization from the literature. Our quantitative and qualitative evaluations highlight the advanced performance of SD-VSum and demonstrate its capacity to generate video summaries that meet the users’ needs. Our contributions are as follows:
@@
  • We extend the recently-introduced VideoXum large-scale dataset for video summarization to S-VideoXum, by producing natural language descriptions of the different human-annotated summaries that are available per video.
@@
  • We propose a new method for script-driven video summarization (SD-VSum) that employs a cross-modal attention mechanism; we use the extended S-VideoXum dataset to train and evaluate this method, and document its competitiveness against other SOTA approaches for query-driven and generic video summarization.
2. Related Work
2.1. Text/Query-driven video summarization
@@
In one of the first attempts towards text-driven video summarization, Sharghi et al. (2016) (Sharghi et al., 2016) proposed a probabilistic model for query-focused video summarization, which selects the video shots that jointly have the highest relevance to the user query and the highest importance in the context of the video. To avoid the use of costly supervision data, in their subsequent work, Sharghi et al. (2017) (Sharghi et al., 2017) described a memory network parameterized probabilistic model that implicitly attends the user query about the video onto different video frames and shots. Vasudevan et al. (2017) (Vasudevan et al., 2017) employed a submodular optimization framework to form a query-adaptive video summarization method that selects the frames of the video summary by quantifying their relevance with the text query based on the distance of deep representations in a common embedding space, and taking into account frames’ diversity, representativeness and aesthetic quality. Wei et al. (2018) (Wei et al., 2018) presented a semantic attended video summarization network that contains a frame selector and a video descriptor, and learns how to choose the most semantically representative video parts by minimizing the distance between the generated description for the video summary and human-provided descriptions of visually-coherent segments of the video. Zhang et al. (2018) (Zhang et al., 2018) used a three-player loss to train a query-conditioned generative adversarial network, where the generator creates a summary based on the learned joint representation of the user query and the video content, and the discriminator tries to discriminate the real summary from a generated and a random one. Jiang et al. (2019) (Jiang and Han, 2019) designed a hierarchical network architecture to model query-related long-range temporal dependency, and combined diverse attention mechanisms used to encode query-related and context-important information, with a multi-level self-attention and a variational autoencoder used to incorporate user-oriented diversity and stochastic factors. Huang et al. (2020) (Huang and Worring, 2020) described a method for query-controllable video summarization, which uses: a summary controller that describes the desired video summary, a summary generator that scores frames based on their relationship with the textual description, and a summary output module that creates a summary by choosing the top-k scoring frames. Xiao et al. (2020) (Xiao et al., 2020b) modeled the high-level semantics of each video shot using local self-attention mechanisms and the semantic relationship between all shots and a given query using a query-aware global-attention, and employed a query-relevance ranking module to compute a similarity score for each shot and a given query, and form the summary. To eliminate the bias when the model is trained on a small dataset, in their next work, Xiao et al. (2020) (Xiao et al., 2020a) trained a hierarchical self-attentive network for estimating the importance of video frames and shots using the ActivityNet Captions dataset (Krishna et al., 2017), fine-tuned it using a reinforced caption generator that constructs a video description using the selected frames, and employed a query-aware scoring module that computes shot-level scores for a given query and produces the video summary. Cizmeciler et al. (2021) (Cizmeciler et al., 2022) presented a query-focused video summarization method that employs semantic attributes as indicators of query relevance and semantic saliency maps from pretrained attribute/action classifiers to locate relevant regions in the video frames. Narasimhan et al. (2021) (Narasimhan et al., 2021) introduced the CLIP-It method for query-focused video summarization, which uses a multi-head language-guided attention mechanism that learns how to estimate the frames importance based on their visual relevance and their correlation with a user-defined query. Formulating the task of query-attentive video summarization as a multi-label classification problem, Hu et al. (2023) (Hu et al., 2023) predicted the correlation between visual content and multi-concept labels by feeding deep features into a multi-layer perceptron, and selected the video part with the highest relevance to the user’s query to form the summary. Finally, to deal with the lack of large labeled datasets for query-based video summarization, Huang et al. (2023) (Huang et al., 2023) resorted to the use of self-supervision and the generation of segment-level pseudo labels from input videos to model the relationship between a pretext and a target task and between the pseudo label and the human-defined label, and employed a semantics booster to generate context-aware query representations and mutual attention to capture the interaction between visual and textual modalities. For further details, we refer the reader to a recent survey on query-attentive video summarization (Kadam and Deshpande, 2024).
@@
The methods described above create a video summary taking into account the user’s preferences about its content, expressed using a few keywords (Sharghi et al., 2016, 2017; Vasudevan et al., 2017; Zhang et al., 2018; Jiang and Han, 2019; Huang and Worring, 2020; Xiao et al., 2020b; Cizmeciler et al., 2022; Xiao et al., 2020a; Narasimhan et al., 2021; Hu et al., 2023; Huang et al., 2023) or a short sentence (Wei et al., 2018; Narasimhan et al., 2021). Contrary to these methods, SD-VSum guides the summarization process using a free-form textual description of the content of the desired video summary, allowing the user to create summaries that are visually and semantically more diverse compared to the summaries generated by the existing text/query-driven summarization methods.
2.2. Multimodal video summarization
@@
Closely related to the text/query-driven videos summarization methods, in terms of the used data modalities, are multimodal approaches for generic video summarization which take as input the full-length video and a textual description of its content and/or the video transcripts. The CLIP-It method of Narasimhan et al. (2021) (Narasimhan et al., 2021) for query-based video summarization is also capable to produce generic video summaries using automatically-generated dense captions of the video content. In their subsequent work, Narasimhan et al. (2022) (Narasimhan et al., 2022) described a transcript-based method for summarizing instructional videos by selecting important steps of the procedure that are most relevant to the task, as well as salient in the video, i.e., referenced in the speech. Zhong et al. (2022) (Zhong et al., 2022) presented an unsupervised video summarization approach that learns how to estimate the frames’ importance and create semantically representative video summaries, by minimizing the distance of learnable representations of the video content and its textual description in a common latent space. He et al. (2023) (He et al., 2023) presented a unified multimodal transformer-based model which can align and attend multimodal inputs leveraging time correspondence (i.e., video and transcript), and selects the most important frames of the video (to form a keyframe-based video summary) and the most importance sentences of the transcript (to produce a text summary). Finally, Argaw et al. (2024) (Argaw et al., 2024) trained a multimodal method for video summarization using ground-truth pseudo-summaries obtained by prompting a Large Language Model to extract the most critical and informative moments from ASR transcripts and mapping back the resulting textual summary to the relevant video segments. Their method creates a video summary by taking into account the visual content and a long-form description of it, or the video transcripts.
@@
2.3. Datasets
Table 1. Overview of the different datasets for text/query-driven video summarization from the literature. The original VideoXum dataset that is used as a basis for our S-VideoXum dataset, is included as a reference.
@@ 2.3. Datasets
UTE (Lee et al., 2012) 1 4 3 - 5 hr. 3 textual summaries train, test
TVE (Yeung et al., 2014) 1 4 45 min. 3 textual summaries train, test
QFVS (Sharghi et al., 2017) 1 4 3 - 5 hr. 46 query-based video summaries train, test
RAD (Vasudevan et al., 2017) 22 190 2 - 3 min. 5 query-relevance labels train, validation, test
SumMe (Gygli et al., 2014) - 25 1 - 6 min. 3-5 fragment-level descriptions train, test
TVSum (Song et al., 2015) 10 50 1 - 11 min. 3-5 fragment-level descriptions train, test
ARS (Cizmeciler et al., 2022) 2 10 2 - 12 min. 10 query-based video summaries train, test
VideoXum (Lin et al., 2024) open domain 14,000 up to 12.5 min. 10 - train, validation, test
@@
An overview of the datasets from the literature for text/query-driven video summarization is given in Table 1. The earliest used ones by Sharghi et al. in (Sharghi et al., 2016), were UT Egocentric (UTE) (Lee et al., 2012) and TV Episodes (TVE) (Yeung et al., 2014). UTE contains four egocentric videos 353-5 hours long that were recorded in an uncontrolled environment from a first-person view wearable camera. TVE includes four videos $45$ min. long that were captured from a third person’s viewpoint. The videos of the UTE and TVE have been annotated by human subjects in (Yeung et al., 2014) with one-sentence descriptions at the level of short video fragments that last $5$ and $10$ sec., respectively. Three reference text-based summaries were defined per video as a subset of the textual annotations, and used for evaluating query-driven summarization based on the ROUGE-SU metric (Lin, 2004). UTE was further extended by Sharghi et al. in (Sharghi et al., 2017), who added human-annotated video summaries for $46$ user queries that were defined using two or three concepts per video, forming the Query-Focused Video Summarization (QFVS) dataset. Vasudevan et al. (Vasudevan et al., 2017) introduced the Relevance and Diversity (RAD) dataset which contains $190$ videos that have been retrieved from YouTube using different queries and annotated with query relevance labels at the frame and fragment level, ranging from 0 (Bad) to 3 (Very Good). This dataset was extended by Huang et al. in (Huang and Worring, 2020), who computed relevance scores for each query-video pair of the dataset. The SumMe (Gygli et al., 2014) and TVSum (Song et al., 2015) benchmarking datasets for generic video summarization were extended by Wei et al. (Wei et al., 2018) with fragment-level descriptions of the visual content, to train a video summarization method that takes into account the video semantics. Finally, Cizmeciler et al. (Cizmeciler et al., 2022) introduced the Activity Related Summaries (ARS) dataset containing $10$ YouTube videos (22-$12$ min. long) that have been annotated by humans at the shot level using a list of $60$ attributes, while $10$ ground-truth summaries were defined per video using a list of $142$ one/two-word queries.
@@
As shown in Table 1, contrary to the existing datasets, the extended S-VideoXum dataset for script-driven video summarization: i) contains a remarkably large number of videos from different domains with diverse visual content, ii) provides $10$ different ground-truth summaries per video, thus allowing to train a summarization method to produce different summaries for a given video based on the users’ preferences about the content of each summary, and iii) is divided into train, validation and test sets, enabling the selection of a well-trained model that generalizes well to unseen data.
3. Problem Statement and Dataset
3.1. Problem formulation